# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v8

@@ 10 Transitive Triangles and Seymour Diamonds
Here, we have $x$ pointing to $y$ and $u$ and $y$ pointing to $u$. The concept of a transitive triangle is an important one because it involves the risk of double counting. The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$. If we do not properly deal with transitive triangles, we might improperly count out second neighbors and thus not have a correct algorithm.
Definition 10.2.
The notion of interior neighbors is another way of describing what [1] referred to as transitive triangles. They showed that these triangles are always present in a graph that doesn’t satisfy Seymour’s conjecture, i.e., one where every node has the decreasing sequence property. We provided a similar result with Lemma 6.2 (Interior Degree at least 1). When a minimum degree node $v_{0}$ has the Decreasing Sequence Property, all of its neighbors have an interior degree of at least 1. That means that every node $x\in N^{+}(v_{0})$ will relate to some other node $y\in N^{+}(v_{0})$ and these three nodes, $(v_{0},x,y)$, will form a transitive triangle. We can extend that further for any two constructive containers.
@@
Lemma 10.4.
Proof.
Suppose we have an oriented graph $G$ with a node $u$ in container $C_{i}$.
@@
11 Main Theorem: Seymour’s Conjecture is True
@@ 11.1 Introduction
We first restate the Lemma 7.5 (Non-Existence of Back-Arcs), from section 6, but slightly rephrased.
Lemma 11.1.
Let $G$ be an oriented graph, with all nodes satisfying the Decreasing Sequence Property (DSP). Let $v_{0}$ be a minimum degree node in $G$. If there are back arcs in $G$, then a node’s degree will double.
@@ 11.2 Theorem Statement
These lemmas are what allowed us to set up and use lemma 6.4 (Weak Decreasing Exteriors Lemma). That lemma gives us a decreasing set of positive numbers. This set is bound to reach the size of two. When it does, one of the nodes in that set will have all of its arcs to the next container (thereby being exterior arcs). This will cause a parent of that node’s degree to double.
@@
1:Initialization:
2:Choose the minimum degree node $v_{0}$.
3:Divide nodes into containers $C_{1},C_{2},\ldots$ based on their degree and proximity to $v_{0}$.
4:for each container $C_{i}$ in $G$ do:
5:     Locate a node $x_{i}$ in $C_{i}$ whose interior degree doubles.
6:     Move to a neighboring container $C_{i+1}$.
7:     if there is a back arc from $C_{i+1}$ to a prior container then
8:         Locate child of $x_{i}$ whose exterior degree doubles, causing $x_{i}$’s overall degree to double.
9:         Halt.
10:     else
11:         Since there is no back arc, the container $C_{i+1}$ will be limited in size to be 1 less than the size of container $C_{i}$.
12:         Repeat from Step 4.
13:     end if
14:end for
@@
In small graphs, such as the one shown in Figure 4.3 (Minimum Out-Degree 3 with Neighbors 1), the collision phenomenon was not observed. This is because the container sizes decrease linearly (at a rate of 1), while the interior degrees increase linearly (also at a rate of 1). However, the number of arcs within the oriented graph grows quadratically, at a rate of $²O(n\texttwosuperior)$. Consequently, while the number of nodes increases linearly, the number of edges increases quadratically. In small graphs, the growth hasn’t progressed far enough for the quadratic edge growth to outpace the linear node growth and cause collisions. In larger graphs, however, this difference in growth rates becomes significant, and collisions begin to occur.
@@
12 Conclusion and Future Work
@@
Future research and development should explore questions of complexity and scalability. Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
Finally, this research offered an alternative method for the SSNC. We focused on identifying a node whose out-degree doubles in the graph’s square rather than just showing that it exists. To do this, we developed an effective algorithm and proved its complexity. Other problems in graph theory and related domains may benefit from this change in emphasis from existential proofs to algorithmic solutions. This may encourage researchers to explore constructive approaches with a focus on practical applications. Hopefully, this work will inspire further research on constructive approaches to other graph theory problems. Also, we encourage greater collaboration between mathematicians and computer scientists to explore the algorithmic and practical implications of theoretical results.
@@ 13 Acknowledgments
I am deeply thankful to God for the inspiration and guidance that shaped this work, as Proverbs 3:6 has been a guiding light throughout this journey. I am profoundly grateful to my family—my mother, father, and wife—for their unwavering support and encouragement.
@@
A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Ranked Structure: The elements of $S$ are partitioned into a ranking of ”containers” or ”levels,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order. The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion, such as a designated metric or property relevant to the application.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
@@ A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
@@ 16 Appendix 3: Visualization and Accessibility of Oriented Graphs
Readers are encouraged to explore the website for additional information and examples at https://glovermethod.com.

### v9

@@ 10.1 Transitive Triangle
Here, we have $x$ pointing to $y$ and $u$ and $y$ pointing to $u$. The concept of a transitive triangle is an important one because it involves the risk of double counting. The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$. If we do not properly deal with transitive triangles, we might improperly count out second neighbors and thus not have a correct algorithm. The notion of interior neighbors is another way of describing what [2] referred to as transitive triangles. They showed that these triangles are always present in a graph that doesn’t satisfy Seymour’s conjecture, i.e., one where every node has the decreasing sequence property. We provided a similar result with Lemma 6.2 (Interior Degree at least 1). When a minimum degree node $v_{0}$ has the Decreasing Sequence Property, all of its neighbors have an interior degree of at least 1. That means that every node $x\in N^{+}(v_{0})$ will relate to some other node $y\in N^{+}(v_{0})$ and these three nodes, $(v_{0},x,y)$, will form a transitive triangle. We can extend that further for any two constructive containers.
@@
Lemma 10.2.
Proof.
Suppose we have an oriented graph $G$ with a node $u$ in container $C_{i}$.
10.2 Seymour Diamond
Definition 10.2.
@@ 10.2 Seymour Diamond
11 Main Theorem
@@ 11.2 Theorem Statement
These lemmas are what allowed us to set up and use lemma 6.4 (Weak Decreasing Exteriors Lemma). That lemma gives us a decreasing set of positive numbers. This set is bound to reach the size of two. When it does, one of the nodes in that set will have all of its arcs to the next container (thereby being exterior arcs). This will cause a parent of that node’s degree to double. ∎
Input: Graph $G$
Output: Halt condition
Initialization:
Choose the minimum degree node $v_{0}$;
Divide nodes into containers $C_{1},C_{2},\ldots$ based on their degree and proximity to $v_{0}$;
for each container $C_{i}$ in $G$ do
       Locate a node $x_{i}$ in $C_{i}$ whose interior degree doubles;
       Move to a neighboring container $C_{i+1}$;
       if there is a back arc from $C_{i+1}$ to a prior container then
             Locate child of $x_{i}$ whose exterior degree doubles, causing $x_{i}$’s overall degree to double;
             Halt;
       end if
      else
             Since there is no back arc, the container $C_{i+1}$ will be limited in size to be 1 less than the size of container $C_{i}$;
             Repeat from Step 4;
       end if
end for
@@
In small graphs, such as the one shown in Figure 4.3 (Minimum Out-Degree 3 with Neighbors 1), the collision phenomenon was not observed. This is because the container sizes decrease linearly (at a rate of O(n)), while the interior degrees increase linearly (also at a rate of O(n)). However, the number of arcs within the oriented graph grows quadratically, at a rate of $O(n^{2})$. Consequently, while the number of nodes increases linearly, the number of edges increases quadratically. In small graphs, the growth hasn’t progressed far enough for the quadratic edge growth to outpace the linear node growth and cause collisions. In larger graphs, however, this difference in growth rates becomes significant, and collisions begin to occur.
@@
12 Applications
The Seymour Second Neighborhood Conjecture is not just a theoretical problem; it has practical applications. We have constructed an algorithm that searches for and finds nodes that satisfy this conjecture. Not only does our algorithm find these nodes, but it partitions the graph into a data structure, which can be beneficial for many other purposes. While the conjecture only asked for a single node, our approach finds a set of degree-doubling nodes, which are critical in understanding both network structure and influence dynamics.
We have developed an algorithm that efficiently identifies degree-doubling nodes. This algorithm has a complexity of $O(|V|+|E|)$, making it scalable to large graphs. In social networks with millions of users, the efficiency of the degree-doubling node identification algorithm is crucial. A less efficient algorithm would make it impractical to analyze the network and identify influential users within a reasonable time-frame.
A/B testing is often considered by companies when thinking of introducing a new product or service to customers. Users are randomly divided into two groups, a control group and a treatment group. It gives data driven decisions that focus on user feedback.
Network A/B testing extends this methodology to social networking experiments, where users are part of an interconnected system. Similar to traditional A/B tests, different versions of a post or an ad are shown to different audiences to compare engagement metrics such as clicks or conversions. However, network A/B testing comes with significant challenges, primarily due to network interference—a phenomenon where users in the control group interact with users in the treatment group, unintentionally influencing the results. This limitation causes some researchers to dismiss network A/B testing entirely.[12]
Network A/B testing often relies on assumptions about influence, which can lead to misleading results. The first assumption we will speak of is homogeneous influence. This says that all users have the same influence, which is clearly not true. However, Influence varies greatly across users, and failing to account for this can distort test results. Assuming all users have the same influence can lead to misrepresenting the results. A small change that affects a highly influential user can appear to have a large overall effect, even if it has little impact on the majority of users. [13]
A second assumption that we will mention is random influence. This assumption says that influence is spread randomly throughout the network. If influence isn’t random, but concentrated in communities, then randomly assigning users to control and treatment groups can lead to imbalances. One group might inadvertently contain a larger proportion of influential users, skewing the results. [14]
Finally, there is the assumption that users are independent of each other and their actions are not influenced by the actions of others. This is an oversimplification, as social networks are built on interconnectedness. This is perhaps the most critical flaw. Users in social networks are not independent. They talk to each other, share information, and influence each other’s opinions. Ignoring these network effects can completely invalidate the results of A/B tests.
Some may make an assumption that only type of influential users are celebrities. These celebrities certainly may have a large name recognition, but they do not represent every influential user. Conversely, every known name is not necessarily an influential user. There are many potential influential users that this assumption misses like community leaders, subject matter experts, connectors between different communities and early adopters.
The Graph Level Order (GLOVER) data structure is a way to organize an oriented graph into leveled containers. Nodes are placed based on their distance from a source node (normally a minimum degree node). Nodes are stored into ordered containers $C_{1}$, $C_{2}$, …, $C_{i}$, where $C_{i}$ contains nodes that are distance $i$ from the source node. Each container then connects to its next container.
The first thing this data structure does is helps with the problem of partitioning data sets. By the definition of the data structure, these sets $C_{i}$ and $C_{i+2}$ have no edges in common. This allows us to more easily set up the control group and the treatment group. The partitioning ensures that there can be no spill-over or network interference between the two groups. The GLOVER data structure, by organizing the graph based on distance from a source node, helps to minimize network interference. Because nodes in different containers (e.g., $C_{i}$ and $C_{i+2}$) have no direct connections, the likelihood of users in the treatment group influencing users in the control group is significantly reduced. This allows for more accurate measurement of the treatment’s true effect.
The GLOVER data structure, combined with our algorithm, allows for parallel processing of the graph. Because the containers in GLOVER are independent, the computation of degree-doubling nodes within each container can be performed concurrently. The ability to parallelize the algorithm significantly reduces the computation time, especially for very large social networks. This makes it feasible to analyze the network and identify influential users in real-time or near real-time, which is essential for dynamic network A/B testing.
Secondly, we have solved the SSNC with a set of degree-doubling nodes. These nodes are are proven to be mathematically influential. Thus, we can go back into the original social network, knowing these nodes are influential with more confidence about both our control and treatment sets. By identifying degree-doubling nodes, we can strategically select users for our A/B test. These nodes are likely to be highly influential within their communities. Including them in the treatment group allows us to observe how the new feature spreads through the network, providing a more realistic assessment of its potential impact. Alternatively, we might choose to exclude them from the treatment group to see how the feature performs without the amplification effect of these key influencers, giving us a different but equally valuable perspective.
Figure 14: Social Media Example shown with and without exterior arcs. The nodes in container 3 are the degree-doubling nodes, aka the influencers. Knowing that, we can develop a more informed treatment group, partitioned from the control group.
While this example uses a simplified network of 30 nodes, our algorithm can efficiently handle much larger networks due to its $O(|V|+|E|)$ complexity and parallelizable nature. In a real-world scenario with millions of users, the computation time would be significantly reduced by parallel processing, making it possible to identify degree-doubling nodes and design network A/B tests efficiently.
This application of the SSNC seeks to plant the seeds for future research in network A/B testing. There are other issues that the GLOVER algorithm does not resolve. Another limitation is that this algorithm will require an oriented graph, instead of a standard directed graph. that may limit applications as well. However, this research may hold a key to unlocking network A/B testing.
13 Conclusion
@@
Future research and development should explore other possibilities for this data structure as well as their complexity and scalability. Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
Finally, this research offered an alternative method for the SSNC, moving beyond existential proofs to a constructive, algorithmic solution. Instead of simply demonstrating the existence of a node whose out-degree doubles in the graph’s square, we focused on identifying such a node through a novel approach. This involved several key innovations.
The first if these techniques is graph partitioning. We dissected the problem by strategically partitioning the graph. This revealed underlying properties that were the about these partitions and how they relate to one another that were crucial to our algorithmic solution. This partitioning allowed us to analyze the relationships between a vertex and its second neighbors in a new light.
Secondly, we employed a divide-and-conquer strategy. After breaking down the problem into smaller, more manageable subproblems, we solved them in an independent nature. Returning these independent subproblems back to the overall problem gave rise to a global solution. This approach, combined with our partitioning technique, enabled us to develop an efficient algorithm.
Lastly, we introduced the concept of dual metrics. This allowed us to simultaneously consider multiple aspects of the problem simultaneously. This multi-faceted perspective proved essential in connecting the local degree properties of vertices within a container to the global neighborhood property of the distance required by the conjecture.
These techniques, particularly the graph partitioning and divide-and-conquer strategies, may be applicable to other problems in graph theory and related domains. While the specific degree and distance dual metrics used in the SSNC proof might not be directly transferable, the general principle of considering multiple perspectives could be valuable. It is possible that other problems will require the development of new data structures and metrics tailored to their specific characteristics. However, the core idea of strategically dissecting the problem, as demonstrated in our solution to the SSNC, offers a promising direction for future research.
This shift in emphasis from existential proofs to algorithmic solutions, combined with the development of new data structures and analytical techniques, may encourage researchers to explore constructive approaches with a focus on practical applications. We hope that this work will inspire further research on such approaches to other graph theory problems. We also encourage greater collaboration between mathematicians and computer scientists to explore the algorithmic and practical implications of theoretical results.
@@ 14 Acknowledgments
I am deeply thankful to God for the inspiration and guidance that shaped this work, as Proverbs 3:5-6 has been a guiding light throughout this journey. I am profoundly grateful to my family—my mother, father, and wife—for their unwavering support and encouragement.
@@
A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
Total Order on Elements of G: We need that every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
@@ A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors if $(x,y)\in G$.
  • If $x\in C_{i}$ and $y\in C_{i+1}$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
@@ 17 Appendix 3: Visualization and Accessibility of Oriented Graphs
Readers are encouraged to explore the website for additional information and examples at
https://glovermethod.com.